OpenStudy (anonymous):

(b) Find the position, velocity, and acceleration of the mass at time t = 5π/6. x(t) = 8 cos t v(t) = -8sint a(t) = -8cost x(5π/6)= v(5π/6)= a(5π/6) =

OpenStudy (anonymous):

i get decimals for the answers, i'm just wonder how i would answer it on a test

OpenStudy (anonymous):

cos(5pi/6) = -sqrt(3)/2 sin(5pi/6) = 0.5

OpenStudy (anonymous):

helps ?

OpenStudy (phi):

you should memorize sin,cos and tan of 0,30,45, 60 and 90 degrees you should know how to change between radians and degrees radians *180/pi or degrees*pi/180 you should know how to figure out the signs depending on which quadrant you are in for sin, cos, tan

OpenStudy (cwrw238):

a = -8 cos (5pi/6)

OpenStudy (anonymous):

okaay so i have to answer in radiants

OpenStudy (phi):

if I were doing this, I would say 5pi/6 is the same as 30 degrees in the 2nd quadrant. sin will be +, cos -

OpenStudy (anonymous):

x(5pi/6)=-4sqroot3

OpenStudy (anonymous):

That is correct for 8cos(5pi / 6)

OpenStudy (anonymous):

okaay

OpenStudy (anonymous):

So you have the concept, just apply it to the rest.

OpenStudy (anonymous):

v(t)=-4

OpenStudy (anonymous):

a(t)=4sqroot3

OpenStudy (anonymous):

v(5π/6)=-4 .. right

OpenStudy (anonymous):

a(5pi/6) right as well

OpenStudy (anonymous):

thank you guys learned alot

OpenStudy (phi):

I don't know that it really helps me memorize sin of these angles, it is interesting: angle sin(angle) 0 \(\frac{\sqrt{0}}{2}\) 30 \(\frac{\sqrt{1}}{2}\) 45 \(\frac{\sqrt{2}}{2}\) 60 \(\frac{\sqrt{3}}{2}\) 90 \(\frac{\sqrt{4}}{2}\)

OpenStudy (anonymous):

thank you